A, B, and C are three complexes of chromium (III) with the empirical formula H12O6CI3Cr. All the three complexes have water and chloride ion as ligands. Complex A does not react with concentrated H2SO4, whereas B and C lose 6.75% and 13.5% of their original mass, respectively, on treatment with concentration H2SO4. Identify A, B and C.

Grade:11

1 Answers

Aditi Chauhan
9 years ago
The complex A does not react with concentrated H2SO4 implying that all water molecules are coordinated with Cr3+ ion. Hence, its structure would be [Cr(H2O)6]CI3.
The compound B loses 6.75% of its original mass when treated with concentrated H2SO4. This loss is due to the removal of water molecules which is/are not directly coordinate to Cr3+ ion.
The mass of water molecules removed per mole of the complex
= 6.75/100 * molar mass f the complex = 6.75/100 * 266.5 g = 17.98 g
This corresponds to one mole of water. Hence, the structure of the compound B will be [Cr(H2O)5CI](H2O)CI2
NOTE : The compound C loses 13.5% of its mass when treated with concentrated H2SO4 which is twice fo the mass lost by the compound B. Hence, the structure of the compound C will be [Cr(H2O)4 CI2](H2O)2 CI.